OpenStudy (anonymous):

Write the equation of a line with slope m = 2 and point (0, 0).

OpenStudy (anonymous):

A. y=0 B. y=2x C. y=2 D. x=2

Vocaloid (vocaloid):

any thoughts so far?

OpenStudy (anonymous):

no

Vocaloid (vocaloid):

plug in your point (0,0) and your slope into the equation y-y1 = m(x-x1) where x1 = 0, y1 = 0, and m = your slope

OpenStudy (anonymous):

ok

Vocaloid (vocaloid):

still there?